OpenStudy (anonymous):

The ordered pair (1,81), (2,100), (3,121), 5,169). represent a function. What is a rule that represent this function?

OpenStudy (anonymous):

The y values are all powers of 6: 6^1, 6^2 = 36, 6^3 = 216, 6^4 = 1296, 6^5 = 7776. x = 1, y = 6^1 x = 2, y = 6^2 x = 3, y = 6^3 x = 5, y = 6^5 So y = 6^x. So the rule is: Raise 6 to the x value to get the y value.

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

is that the answewr

OpenStudy (anonymous):

@elementwielder

OpenStudy (aum):

The y-values are all perfect squares. x y 1 81 = 9^2 2 100 = 10^2 3 121 = 11^2 5 169 = 13^2 When x is 1 we square 9 to get the y value (that is, 8 is added to x and squared to get y) When x is 2 we square 10 to get the y value (that is, 8 is added to x and squared to get y) Therefore, y = (x+8)^2 You can try putting various x values such as 1,2,3,5 and see if it gives the y-value in the table to make sure y = (x+8)^2 is the correct answer.

OpenStudy (anonymous):

yes thas wat i have thank you very much

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

XD